OpenStudy (anonymous):

need help please... Write the point-slope form of the equation passing through (5, -1) with a slope of 6. A. y-1=6(x+5) B. y+5=6(x-1) C. y+1=6(x-5) D. y-5=6(x+1)

jimthompson5910 (jim_thompson5910):

Start with y - y1 = m(x - x1) which is point slope form and plug in m = 6 (the slope) and (x1, y1) = (5, -1), which is the given point

OpenStudy (anonymous):

so would be y-1=6(x-5)?

OpenStudy (anonymous):

You have a -1 for your y so you have to subtract a negative number, which is adding a positive number.

OpenStudy (anonymous):

y-(-1)=6(x-5) and then simplify the left side.

OpenStudy (anonymous):

so would it be a?

jimthompson5910 (jim_thompson5910):

keep in mind that subtracting a negative is like adding ex: 2-(-3) = 2+3 11-(-7) = 11+7 x-(-z) = x+z

OpenStudy (anonymous):

ok so it would be y+5 correct?

OpenStudy (anonymous):

i mean y+1 ?

jimthompson5910 (jim_thompson5910):

yes y-(-1) turns into y+1

OpenStudy (anonymous):

thank you for helping me! :)

jimthompson5910 (jim_thompson5910):

np